The following is a known probability model for bills collected at a bank.

x | $1 | $2 | $5 | $10 | $20 | $50 | $100

P(x) | 0.35 | 0.01 | 0.25 | 0.2 | 0.1 | 0.05 | 0.04

-What is the probability that the first bill of the day that is collected is a $20 bill, the second bill collected is a $1 bill, and the third bill is a $10 bill?

-What is the probability that the first bill is not a $2 bill, the second bill is not a $2 bill and the third is not a $2 bill?
Answer #1
X P(X)
1 0.35
2 0.01
5 0.25
10 0.2
20 0.1
50 0.05
100 0.040

probability that the first bill of the day that is collected is a $20 bill, the second bill collected is a $1 bill, and the third bill is a $10 bill = 0.1*0.35*0.2=0.007
---------------

probability that the first bill is not a $2 bill, the second bill is not a $2 bill and the third is not a $2 bill = 0.99*0.99*0.99 = 0.970299